[Brahms-l] Shift report 20050607 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 07 2005 - 21:15:18 EDT
Shifter leader was BRK
Shifters were 


The measured currents for D2 are not within the acceptable ranges.
Set point 0.00 and readout value at 2936.28.

After the dump, the runs have been ended.

The runs were from 15305 to 15310.

Beam dumped. 

Experiments ended at about 8.30 PM.

D2 magnets is showing strange behaviour when ramped down. There is residual field. 
The message has been given above.

The crucial figures were not identical to the standard copies shown on the wall.

Leaving early, as instructed by JH before the due hour of 24.00


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 7 21:15:31 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 07 2005 - 21:15:44 EDT